A function f(x) satisfies f(x+y)=f(x)+y for all x, y belongs to R and f(0)=5. Then f(2020) is equal to 

Grade:12

1 Answers

divyam gupta
5 years ago
f(x+y)=f(x)+y ...(given)
let’s take x=0,y=1
f(0)=5....(given)
according to condition
f(0+1)=f(0)+1=6
put, x=0,y=2
f(0+2)=f(0)+2=7
similarly,put x=0and y=2020
f(0+2020)=f(2020)=f(0)+2020=2025
 
ans=2025